The length of a rectangle is 2cm more than twice its width. If the perimeter of the rectangle is 52cm, find the dimensions of the rectangle.
Answer by Nate(3500)

Length = 2w + 2
Width = w
Equation:
2w + 2l = 52
2w + 2(2w + 2) = 52
2w + 4w + 4 = 52
6w = 48
w = 8
The width is eight units, and the length is 18 units.